Question – 5: Consider the vectors,a=12i + 2j, b= -2i + 3j, and c = a-b
(a) Find the magnitude, and the direction of the vectors a,b and c.

(b) Find the value of the angle between a and b .

(c) Find the area of the triangle whose vertices are the end points of the vectors, a , b , and c .
: Question – 5: Consider the vectors,a=12i + 2j, b= -2i + 3j, and c = a-b
(a) Find the magnitude, and the direction of the vectors a,b and c.

(b) Find the value of the angle between a and b .

(c) Find the area of the triangle whose vertices are the end points of the vectors, a , b , and c .

Answer by stanbon(19009)

Consider the vectors,a=12i + 2j, b= -2i + 3j, and c = a-b
(a) Find the magnitude, and the direction of the vectors a,b and c.
Looking at vector "a" = 12i+2j
magnitude = sqrt(12^2+2^2) = sqrt148 = 2sqrt37
direction = arc tan (2/12) = 9.4623 degrees
Comment: Do the same for b and c
----------------------
(b) Find the value of the angle between a and b .
Look at your direction angles for and b and take the appropriate difference.
-------------------
(c) Find the area of the triangle whose vertices are the end points of the vectors, a , b , and c .
This depends alot on the formulas you have for the area of a triangle
-------------
Cheers,
Stan H.